The test statistic of z=2.51 is obtained when testing the claim that p≠0.649

a. Identify the hypothesis test as being​ two-tailed, left-tailed, or​ right-tailed.

b. Find the​ P-value.

c. Using a significance level of

alphaαequals=0.100.10​,

should we reject

Upper H 0H0

or should we fail to reject

Upper H 0H0​?
Answer #1

This is Two-tailed because of inequality \tiny \ne .

P-value = 2*P(z>2.51) = 2*0.00604 = 0.0121 (rounded to 4 decimal places)

p-value is less than alpha hence we Reject H0.

Please comment if any doubt. Thank you.